OpenStudy (rebeccapink6398):

Find the discount. 1.regular price=420$; discount rate= 1/4 _______________________________________ 2.regular price=168; discount rate=30% _______________________________________ Ok so I have a question... Do I just - the percent to the number? OHH!!!! yay!!! @31356 ur here :D

OpenStudy (rebeccapink6398):

:D :D :D

OpenStudy (31356):

:D:D:D

OpenStudy (31356):

@texaschic101 can help us too! :D

OpenStudy (rebeccapink6398):

Yes :D she was just helping me :D

OpenStudy (31356):

Oh lol, :D

OpenStudy (rebeccapink6398):

Find the discount. 1.regular price=420$; discount rate= 1/4 _______________________________________ 2.regular price=168; discount rate=30% _______________________________________ Ok so I have a question... Do I just - the percent to the number? :D

OpenStudy (31356):

Discount = 420*1/4 (FOR THE FIRST ONE)

OpenStudy (31356):

I got 105 for the first one?

OpenStudy (rebeccapink6398):

:D

OpenStudy (texaschic101):

.25(420) = 105....correct

OpenStudy (31356):

Thanks for the medal! :D

OpenStudy (texaschic101):

second one is 0.30(168) = 50.40

OpenStudy (31356):

Yea that's what I got too, lol :D

OpenStudy (texaschic101):

wait...are you looking for the discount or the price after the discount is applied ?

OpenStudy (31356):

Right......

OpenStudy (rebeccapink6398):

question 1...done questions 2...done Question 3. the regular price of an item is 35$. The item is on sale with a discount of 10$ what is the discount rate? question 4. what is the discount rate for an $80 music player that is on sale for 68$? :D

OpenStudy (31356):

For the third one discount of $10 or 10%?

OpenStudy (rebeccapink6398):

$10

OpenStudy (31356):

Oh okay...

OpenStudy (31356):

Sorry, got to go.........

OpenStudy (rebeccapink6398):

helo @texaschic101 ? :)

OpenStudy (rebeccapink6398):

help*

OpenStudy (texaschic101):

3. 10/35 = 0.285 = 28.5% ---I am not sure about this one 4. 80 - 68 = 12 : 12/80 = 0.15 = 15%

OpenStudy (rebeccapink6398):

hmmm..... im not sure either :(